# Adaptive Layer-skipping in Pre-trained LLMs

### v2

@@ 3.3 FlexiDepth on Different Language Models
We evaluate FlexiDepth across language models of varying sizes to assess its adaptability. All models, which are instruction-tuned, were trained using the same configuration and dataset as outlined in our implementation details. Larger models, namely Llama-2-13B-Instruct and Llama-3-8B-Instruct, skip more layers compared to the smaller Qwen-2.5-3B-Instruct [39]. Specifically, Llama-2-13B-Instruct and Llama-3-8B-Instruct skip an average of 6.42 and 6.25 layers, respectively, across benchmarks, while Qwen-2.5-3B-Instruct skips only about 1.49 layers. This trend indicates that larger models possess greater inherent redundancy, enabling more aggressive layer skipping without notable performance degradation. Given this trend, FlexiDepth can potentially be applied to even larger models to further exploit their redundancy. For instance, it could be applied to large mixture-of-experts models like DeepSeek-V3 [7], where experts are distributed across multiple servers during inference. By skipping entire layers, FlexiDepth could reduce cross-server communication overhead, enhancing inference efficiency of MoE models.
Model MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Llama-2-13B 0.492 0.727 0.714 0.236 0.102 0.790 100.0%
FlexiDepth 0.481 0.741 0.716 0.253 0.096 0.780 100.2%
[…]
Llama-3-8B 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiDepth 0.663 0.743 0.756 0.657 0.323 0.803 102.1%
[…]
Qwen-2.5-3B 0.651 0.702 0.639 0.576 0.487 0.705 100.0%
FlexiDepth 0.643 0.701 0.679 0.583 0.476 0.741 101.5%
[…]
@@ 3.4 Layer allocation dataset
To investigate these layer-skipping patterns and their relationship to task complexity, we constructed a dataset using FlexiDepth with Llama-3-8B-Instruct as the base model. This dataset captures layer usage for token generation in two key domains: language comprehension and math reasoning. By recording the number of layers utilized per token, we aim to uncover how FlexiDepth dynamically adjusts its depth to meet varying task demands.
@@
For text generation tasks, we constructed a dataset by randomly sampling 100 paragraphs from the XSum test set [28], which is a collection of news articles. We evaluated FlexiDepth on three subtasks for each paragraph: copying, summarization, and continuation:
@@
The layer usage pattern across these three tasks was as follows: repeating averaged 20.09 layers (variance 8.08), addition averaged 22.45 layers (variance 14.67), and multiplication averaged 23.90 layers (variance 21.61). These findings align with task complexity—repetition, a straightforward memory task, uses the fewest layers, while multiplication, requiring iterative computation, demands the most. The higher variance in multiplication (21.61) compared to addition (14.67) and repetition (8.08) indicates that computational tokens exhibit more diverse layer needs, likely due to the increasing complexity of intermediate steps. Appendix A provides additional results, showing that tokens tied to arithmetic operations consistently require more layers than directly copying numbers from the context.
@@
3.5 Layer utilization pattern
To complement the token-centric analysis in prior subsections, we examine the layer utilization of FlexiDepth on Llama-3-8B-Instruct. We measure the percentage of tokens processed by each transformer layer across tasks outlined earlier: text generation (copying, summarization, continuation) and math reasoning (repeating, addition, product). Figure 4 visualizes the results, with the left plot showing language tasks and the right plot showing math reasoning. Layer utilization displays a bowl-like pattern, with higher token processing at the initial and final layers, and reduced usage in the middle layers. For language tasks, copying exhibits the deepest dip in middle-layer usage, while continuation involves the most middle layers, reflecting varying task complexity. Similarly, in math reasoning, repeating shows the least middle-layer usage, whereas product requires the most, consistent with increasing task difficulty. The overall layer usage distribution follows a long-tail pattern, where most tokens are processed by a few layers, and others are selectively skipped. We hypothesize that early layers primarily parse and understand input context, middle layers perform task-specific processing that varies with complexity, and final layers focus on decoding for coherent output generation. This functional specialization explains the bowl-like pattern: early and final layers are essential for nearly all tokens, while middle layers are skipped more frequently for simpler tasks that require less intermediate processing. While this explanation remains a hypothesis, future studies are needed to verify the specialized roles of different layers.
Model MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Llama-3-8B 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiExit 0.676 0.741 0.756 0.681 0.220 0.787 96.9%
[…]
Table 3: Performance and skipping patterns of FlexiExit
3.6 Early exit
We extend FlexiDepth’s layer-skipping mechanism into an early-exit framework called FlexiExit, which allows tokens to exit at intermediate layers. In FlexiExit, we adapt FlexiDepth’s router and adaptor modules so that a token flagged for skipping at a given layer bypasses the attention modules in all subsequent layers and is processed solely by lightweight adaptors. We applied FlexiExit to the Llama 3 8B Instruct model, training it under the same conditions as FlexiDepth but with a reduced layer-skipping weight of $\alpha=3\times 10^{-4}$. As shown in Table 3, FlexiExit retains the performance of the original Llama-3-8B-Instruct, while skipping an average of 1.61 to 4.68 layers. The current design of the router and adapter remains effective in supporting early exit scenarios.
@@ 3.6 Early exit
3.7 Ablation studies
We show our ablation studies in this part. All experiments in this section utilize the same layer-skipping penalty, dataset, and training parameters as the original FlexiDepth implementation on the pre-trained Llama-3-8B-Instruct.
@@ 3.7 Ablation studies
In FlexiDepth, we employed a bottlenecked MLP as the router to facilitate nuanced routing decisions. When replacing this component with a simpler linear layer followed by a sigmoid activation, as previously used in mixture-of-depth (MoD) [32], we observe a substantial decrease in performance, particularly on the math reasoning task. Our analysis shows that the linear router mistakenly assigns too few layers to tokens requiring complex computations, significantly reducing the GSM8K performance from 0.657 to 0.131. Overall, the linear router retains only 78.6% of the original performance, emphasizing the importance of the bottlenecked MLP for routing.
@@
We conducted an ablation study in which the KV cache is not computed for tokens routed to skip layers (illustrated as "No KV Cache" in Figure 3). Removing the KV cache results in substantial performance decline, retaining only 87.4% of the baseline performance. This highlights the need to retain contextual information during generation.
@@
Many layer skipping methods do not have adapters. We therefore investigate the performance of FlexiDepth without adapter. This simplification drastically reduces model performance to merely 32.2% of the original FlexiDepth results. The adapter proves critical in aligning latent representations between skipped and processed states, ensuring representation coherence across the model’s layers.
@@ 4 Limitations
Although FlexiDepth reduces FLOPs by adaptively skipping transformer layers, our implementation does not lead to improved throughput on the existing GPU hardware. In FlexiDepth, samples within the same batch can take different execution paths during decoding. Consequently, each FlexiDepth layer must handle tokens that undergo full processing alongside those that skip layers, requiring simultaneous management of both computation and skip branches. This introduces overhead from control-flow management and irregular memory accesses, outweighing the speedup of theoretical FLOP reductions. Future work could investigate hardware-oriented optimizations to address these bottlenecks, leveraging techniques such as token grouping [31], expert sharding [31], and load balancing [16] to better align with GPU architectures. Such hardware-tailored enhancements could further unlock the efficiency potential of FlexiDepth in practice.
@@ 5 Related works
The existing language models can be broadly classified into three categories: encoder-only models (e.g., BERT [8]), encoder-decoder models (e.g., T5 [30]), and decoder-only models (e.g., GPT [3]). Across these architectures, various conditional computation methods have been proposed to optimize resource allocation by dynamically processing tokens.
In encoder-only models, conditional token processing has been explored to prune less critical tokens during processing. For instance, PoWER-BERT [12] and LTP [20] utilize attention scores to assess token importance, dynamically eliminating those deemed less significant. Similarly, LoT [21] introduces a router at each layer to determine whether tokens should be processed or skipped, tailoring computation to token relevance. These approaches leverage the bidirectional nature of encoders, allowing noncausal routing decisions based on full sequence context.
For encoder-decoder models, methods such as CoDA [22] and COLT5 [1] enable conditional computation within the encoder component. By replacing standard attention or feed-forward network (FFN) modules with lightweight alternatives for certain tokens, these techniques assign varying computational resources based on token complexity. However, their noncausal design constrains their applicability to decoder.
In decoder-only models, the autoregressive nature poses unique challenges for dynamic token processing, as tokens must be generated sequentially while preserving contextual dependencies. Piorneering efforts like MoD [32] and SkipLayer [42] address this by deploying a router at each Transformer layer, enabling tokens to dynamically bypass it. These models are trained end-to-end from scratch, integrating the routing mechanism into the learning process. Duo-LLM [2] also employs a router for conditional computation, but with a distinct method. It first constructs a dataset of oracle routing strategies via exhaustive enumeration and then trains the router on this data. While effective, these methods demand significant computational resources for training and dataset preparation, making them impractical for adapting pre-trained models. MindSkip [13] enables skipping in pre-trained models via learnable routers. It finds that bypassing FFN modules or entire layers significantly degrades performance and instead proposes a method to skip the attention module efficiently. Unlike these approaches, our method, FlexiDepth, enables adaptive layer-skipping in pre-trained models without compromising generation performance by using a plug-in router and adapter architecture.
@@ 6 Conclusion
We present FlexiDepth, an approach to dynamically skip layers in pre-trained large language models. It achieves state-of-the-art performance without altering the original model parameters. FlexiDepth offers valuable insights into the varying computational demands of token generation. We constructed a dataset capturing these allocation patterns to encourage further study of computational needs of tokens, thereby advancing the development of more adaptive and resource-efficient language models. Future works can be done for the application of FlexiDepth in practice..
7 Acknowledgements
This work was partially supported by the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487.
@@ Appendix A Layer allocation Examples
In this section, we present examples to illustrate the layer allocation patterns of FlexiDepth, using the pre-trained Llama-3-8B-Instruct model [9] as the base.
@@ A.1 Language Task
Please copy this paragraph: $<$paragraph$>$Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.$<$/paragraph$>$ Directly output the copied paragraph here:
@@
Please summarize this paragraph into a single sentence: $<$paragraph$>$Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.$<$/paragraph$>$ Directly output the summarized paragraph here:
@@
Please continue writing this paragraph: $<$paragraph$>$Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.$<$/paragraph$>$ Directly output the continued paragraph here:

### v3

@@ 3.3 FlexiDepth on different language models
We evaluate FlexiDepth on language models of varying sizes. These models are trained under the identical configuration and dataset described in our implementation details. Larger models, specifically Llama-2-13B-Instruct and Llama-3-8B-Instruct, exhibit greater number of layers skipped compared to the smaller Qwen-2.5-3B-Instruct (Yang et al., 2024a). Specifically, Llama-2-13B-Instruct and Llama-3-8B-Instruct skip an average of 6.42 and 6.25 layers, respectively, across benchmarks, while Qwen-2.5-3B-Instruct skips only about 1.49 layers. This trend indicates that larger models possess greater inherent redundancy, enabling more aggressive layer skipping without notable performance degradation. Given this trend, FlexiDepth can potentially be applied to even larger models to further exploit their redundancy. For instance, it could be applied to large mixture-of-experts models like DeepSeek-V3 (DeepSeek-AI et al., 2024), where experts are distributed across multiple servers during inference. By skipping entire layers, FlexiDepth could reduce cross-server communication overhead, enhancing inference efficiency of MoE models.
@@ 3.4 Layer allocation dataset
To investigate these layer-skipping patterns and their relationship to task complexity, we constructed a dataset using FlexiDepth with Llama-3-8B-Instruct as the base model. This dataset captures layer usage for token generation in two key domains: language modeling and math reasoning. By recording the number of layers utilized per token, we aim to uncover how FlexiDepth dynamically adjusts the layers to meet varying task demands.
@@
For text generation tasks, we constructed a dataset by randomly sampling 100 paragraphs from the XSum test set (Narayan et al., 2018), which is a collection of news articles. We evaluated FlexiDepth on three subtasks for each paragraph: copying, summarization, and continuation:
@@
The layer usage pattern across these three tasks was as follows: repeating averaged 20.09 layers (variance 8.08), addition averaged 22.45 layers (variance 14.67), and multiplication averaged 23.90 layers (variance 21.61). Appendix A provides additional results, showing that tokens tied to arithmetic operations consistently require more layers than directly copying numbers from the context.
3.5 Layer utilization pattern
@@
To complement the token-centric analysis in prior subsections, we examine the overall layer utilization pattern of FlexiDepth on Llama-3-8B-Instruct. We measure the percentage of tokens processed by each transformer layer across the tasks outlined earlier: text generation (copying, summarization, continuation) and math reasoning (repeating, addition, product). Figure 4 visualizes the results, with the left plot showing language modeling tasks and the right plot showing math reasoning tasks. Layer utilization exhibits a bowl-like pattern, with higher token processing at the initial and final layers, and reduced usage in the middle layers. For language tasks, copying exhibits the deepest dip in middle-layer usage, while continuation utilizes the most middle layers, reflecting varying task complexity. Similarly, in math reasoning, repeating shows the least middle-layer usage, whereas product requires the most, which is consistent with increasing task difficulty. The overall layer usage distribution approximately follows a long-tail pattern, where most tokens are processed by a few layers, and others are selectively skipped. We hypothesize that early layers primarily parse and understand input context, middle layers perform task-specific processing that varies with complexity, and final layers focus on decoding for coherent output generation. Future studies are needed to verify the specialized roles of different layers.
3.6 Ablation Studies
We present ablation studies in this section. All models are based on Llama-3-8B-Instruct and trained using identical configurations as described in our implementation details.
@@ 3.6 Ablation Studies
MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Vanilla 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
$\alpha=5e-4$ 0.671 0.730 0.740 0.687 0.347 0.806 103.7%
Skipped 2.32 0.98 1.03 3.49 6.40 4.01 -
$\alpha=1e-3$ 0.663 0.743 0.756 0.657 0.323 0.803 102.1%
Skipped 4.12 2.00 3.97 10.42 9.55 7.44 -
$\alpha=2e-3$ 0.637 0.721 0.736 0.528 0.140 0.778 86.6%
Skipped 6.10 3.85 5.60 12.53 11.88 9.92 -
Table 4: The influence of the coefficient $\alpha$ for layer skipping.
@@
In FlexiDepth, we employed a bottlenecked MLP as the router to facilitate nuanced routing decisions. When replacing this component with a simpler linear layer followed by a sigmoid activation, as previously used in mixture-of-depth (MoD) (Raposo et al., 2024), we observe a substantial decrease in performance, particularly on the math reasoning task. Our analysis shows that the linear router mistakenly assigns too few layers to tokens requiring complex computations, significantly reducing the GSM8K performance from 0.657 to 0.131. Overall, the linear router retains only 68.7% of the original performance, emphasizing the importance of the bottlenecked MLP for routing.
@@
We conducted an ablation study in which the KV cache is not computed for tokens routed to skip layers (illustrated as "No KV Cache" in Figure 3). Removing the KV cache results in substantial performance decline, retaining only 84.3% of the original performance. This highlights the need to retain contextual information during generation.
@@
Many layer skipping methods do not have adapters. We therefore investigate the performance of FlexiDepth without adapter. This simplification drastically reduces model performance to merely 28.1% of the original FlexiDepth results. The adapter proves critical in aligning latent representations between skipped and processed states, ensuring representation coherence across the model’s layers.
Different Coefficient.
We investigate the impact of the coefficient $\alpha$ in the layer skipping loss (Equation 5). With $\alpha=5\mathrm{e}{-4}$, the model only skips an average of 3.04 layers. Increasing to $\alpha=2\mathrm{e}{-3}$ results in more layers being skipped (averaging 8.31 layers) but causes significant performance drops on reasoning tasks, with GSM8K decreasing to 0.528 and HumanEval to 0.140. The default value $\alpha=1\mathrm{e}{-3}$ achieves a better balance.
@@ 4 Limitations
Although FlexiDepth reduces FLOPs by adaptively skipping transformer layers, our implementation does not lead to improved throughput on the existing GPU hardware (see Appendix B). In FlexiDepth, samples within the same batch can take different execution paths during decoding. Consequently, each FlexiDepth layer must handle tokens that undergo full processing alongside those that skip layers, requiring simultaneous management of both computation and skip branches. This introduces overhead from control-flow management and irregular memory accesses, outweighing the speedup of theoretical FLOP reductions. We are currently exploring approaches to achieve actual throughput improvements with FlexiDepth and have found that its framework can be adapted for early exit strategies (see Appendix C). Future work could also investigate hardware-oriented optimizations to address these bottlenecks, leveraging techniques such as token grouping (Rajbhandari et al., 2022), expert sharding (Rajbhandari et al., 2022), and load balancing (Huang et al., 2024) to better align with GPU architectures. Such hardware-tailored enhancements could further unlock the efficiency potential of FlexiDepth in practice.
@@ 5 Related works
The existing language models can be broadly classified into three categories: encoder-only models (e.g., BERT (Devlin et al., 2019)), encoder-decoder models (e.g., T5 (Raffel et al., 2020)), and decoder-only models (e.g., GPT (Brown et al., 2020)). Across these architectures, various conditional computation methods have been proposed to optimize resource allocation by dynamically processing tokens.
In encoder-only models, conditional token processing has been explored to prune less critical tokens during processing. For instance, PoWER-BERT (Goyal et al., 2020) and LTP Kim et al. (2022) utilize attention scores to assess token importance, dynamically eliminating those deemed less significant. Similarly, LoT (Kim et al., 2023) introduces a router at each layer to determine whether tokens should be processed or skipped, tailoring computation to token relevance. These approaches leverage the bidirectional nature of encoders, allowing noncausal routing decisions based on full sequence context.
For encoder-decoder models, methods such as CoDA (Lei et al., 2023) and COLT5 (Ainslie et al., 2023) enable conditional computation within the encoder component. By replacing standard attention or feed-forward network (FFN) modules with lightweight alternatives for certain tokens, these techniques assign varying computational resources based on token complexity. However, their noncausal design constrains their applicability to decoder.
In decoder-only models, the autoregressive nature poses unique challenges for dynamic token processing, as tokens must be generated sequentially while preserving contextual dependencies. Piorneering efforts like MoD (Raposo et al., 2024) and SkipLayer (Zeng et al., 2023) address this by deploying a router at each Transformer layer, enabling tokens to dynamically bypass it. These models are trained end-to-end from scratch, integrating the routing mechanism into the learning process. Duo-LLM (Alizadeh et al., 2024) also employs a router for conditional computation, but with a distinct method. It first constructs a dataset of oracle routing strategies via exhaustive enumeration and then trains the router on this data. MindSkip (He et al., 2024a) enables skipping in pre-trained models via learnable routers. It finds that bypassing FFN modules or entire layers significantly degrades performance and instead proposes a method to skip the attention module efficiently. Unlike these approaches, our method, FlexiDepth, enables adaptive layer-skipping in pre-trained models without compromising generation performance by using a plug-in router and adapter architecture.
@@ 6 Conclusion
We present FlexiDepth, an approach to dynamically skip layers in pre-trained large language models. It achieves state-of-the-art performance without altering the original model parameters. FlexiDepth offers valuable insights into the varying computational demands of token generation. We constructed a dataset capturing these allocation patterns to encourage further study of computational needs of tokens, thereby advancing the development of more adaptive and resource-efficient language models.
Acknowledgements
This work was partially supported by the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487. We would like to thank Facebook (now Meta) for donating the A100-40G GPUs for conducting the experiments.
@@ Appendix A Layer allocation Examples
In this section, we present examples to illustrate the layer allocation patterns of FlexiDepth, using the pre-trained Llama-3-8B-Instruct model (Dubey et al., 2024) as the base.
@@ A.1 Language Task
Please copy this paragraph: <<paragraph>>Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.<</paragraph>> Directly output the copied paragraph here:
@@
Please summarize this paragraph into a single sentence: <<paragraph>>Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.<</paragraph>> Directly output the summarized paragraph here:
@@
Please continue writing this paragraph: <<paragraph>>Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.<</paragraph>> Directly output the continued paragraph here:
@@ A.2 Math Task
Appendix B Throughput
We evaluated the throughput of FlexiDepth compared to the vanilla Llama-3-8B-Instruct (Dubey et al., 2024) across the paper’s benchmarks. Experiments were conducted on 8 A6000 GPUs with a batch size of 8 and 5-shot examples. For multi-token generation tasks, we fixed the output length at 5 tokens to ensure consistent comparison. All measurements were obtained using the lm-evaluation-harness toolkit (Gao et al., 2024), with throughput reported as iterations per second. Table 5 presents the results across six benchmarks.
Method MMLU Hellaswag Winogrande GSM8K HumanEval CoQA
Vanilla 32.21 11.19 26.55 5.53 10.34 8.29
FlexiDepth 33.43 11.20 26.17 5.46 10.18 7.78
Table 5: Throughput (iterations/s) comparison across benchmarks.
Appendix C FlexiExit
We extend FlexiDepth’s layer-skipping mechanism into an early-exit framework called FlexiExit, which allows tokens to exit at intermediate layers. In FlexiExit, we adapt FlexiDepth’s router and adaptor modules so that a token flagged for skipping at a given layer bypasses the attention modules in all subsequent layers and is processed solely by lightweight adaptors. We applied FlexiExit to the Llama-3-8B-Instruct model, training it under the same conditions as FlexiDepth but with a reduced layer-skipping weight of $\alpha=3\times 10^{-4}$. As shown in Table 6, FlexiExit retains the performance of the original Llama-3-8B-Instruct, while skipping an average of 1.61 to 4.68 layers. The current design of the router and adapter remains effective in supporting early exit scenarios.
Model MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Llama-3-8B 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiExit 0.676 0.741 0.756 0.681 0.220 0.787 96.9%
Skipped 2.17 1.81 3.63 4.68 1.61 3.16 -
Table 6: Performance and skipping patterns of FlexiExit